OpenStudy (goldphenoix):

What does 8th grade math covered? I want to learn about it earlier, so I can become pro in math when I get to 8th grade. :)

OpenStudy (anonymous):

largely depends on the school elementary algebra maybe?

OpenStudy (zarkon):

Banach Spaces

OpenStudy (anonymous):

your best bet is to make sure you have a very good foundation in arithmetic, and elementary algebra you don't want to be spending time trying to figure out what \(\frac{2}{5}-\frac{3}{4}\) is

OpenStudy (goldphenoix):

I learned that. o_O I learned a lot of things. Right now, I'm learning binomial.

mathslover (mathslover):

8th grade math usually covers : Rational Numbers Linear equations in one Variable Understanding quadrilaterals Practical geometry Data Handling Squares and square roots Cubes and cube roots comparing quantities Visualizing in solid shapes Algebraic expressions Mensuration Exponents and Powers Direct and Inverse proportion Factorization Introduction to graphs Playing with numbers

OpenStudy (goldphenoix):

Interesting. I already know more than half of it. :O

mathslover (mathslover):

Good! That was of CBSE Board - India . Check this : ICSE Board - India : http://icse.meritnation.com/icse-math/real-numbers/631/3622/tutorials.html

OpenStudy (goldphenoix):

Interesting website. You all deserve medal, but I'm going to give it to mathslover (He doesn't have a medal, yet. :O)

OpenStudy (anonymous):

are you taking algebra

OpenStudy (goldphenoix):

Yeah. I'm up to Binomial.

OpenStudy (goldphenoix):

Well, I don't kow the difference between pre-algebra and algebra.

mathslover (mathslover):

Thanks for the medal. Best of luck, start learning in advanced. That will help you a lot.

OpenStudy (anonymous):

well in 8th grade u can take algebra or regular math i took algebra

mathslover (mathslover):

Pre-algebra essentially introduces the concept of the single variable, where one part of a binary operation (addition, subtraction, multiplication, or division) is replaced by a box. The box, in algebra, is replaced by a letter such as x or y, to denote a variable, or a, b, or c to denote constants. For example: X + 5 = 9 Pre-algebra also incorporates fractions, decimals, conversions between them, and the standard mathematical operations. In short, everything you need to build upon the concepts for algebra. Algebra begins with single variable equations and builds up to systems of equations with multiple variables, quadratic equations, polynomials (equations of degree n), real and irrational numbers, the cartesian coordinate system, and on up to matrix operations involving multiple dimensions. I believe algebra pretty much stops short of Hilbert space, which has an infinite number of dimensions. ax + by = c is a typical line equation from algebra which, in slope intercept form, is: y = mx + b. b = the y-intercept, and m=∆y/∆x or (y2 - y1)/(x2 - x1) ax2 + bx + c = d is a typical quadratic equation. Polynomials would be of the form axn + bx(n-1) + ... + z = 0. Factoring polynomials is another important concept from algebra. If you can do long division by hand you have some of the key concepts for factoring already. Perhaps the most useful concept from algebra is ratios and proportions. - Source : http://wiki.answers.com/Q/What_is_the_difference_between_pre-algebra_and_algebra_1

OpenStudy (anonymous):

i hated proportions but algebra is really easy if you catch on quick i got a 98 on the regents